SW&IT Meeting Minutes 2020-06-22
Attendees: Carl, David, Igor, Matthieu, Stephen, Xianyu, Yang
Report auto-generated every Monday at 6:30am, Do not edit this table, Update your entries in smartsheets.
| LR20200615 - Bec, Matthieu
| LR20200615 - Holmberg, Carl
| LR20200615 - Edmeades, David
| LR20200615 - Sola, Igor
| LR20200615 - Hooper, Stephen
| LR20200615 - Zhang, Yang
| LR20200615 - Zhang, Xianyu
News from management
The current LBTO COVID-19 Safety Protocols for "Mountain" (M004s00753) and "Tucson" (M004s00752) locations are now posted in the "Information for All Staff" "COVID Related Information" folder, and attached:
Please read these documents!
- tighten our control with operation machines (anacondo install on luci?)
Previous Action Items
- GCSGUI debugging and patching (seems to be working fine now).
- TMS meeting and software work.
- AO-Cravat meeting and wiki page work.
- TMS software work (hopefully some updates at the TMS meeting).
- TMS meeting on Thursday.
- AO-Cravat work (software, wiki etc.).
- Safety training.
- Last week
- AO/SOUL operation support
- UAO skip frame SW change, test
- SOUL skip frame SW test and support
- Add record paramselector in OptLoop Diagnostic SW change and test
- This week
- SOUL operation support
- Paramselector SW/configration change and test
- Power Backplane action items
- Laura new computer
- Working with Dell rep since it's so new
- ML new computer
- Gene says there's a problem with the tradin and LBTC rules, MB to follow up in OC
- JC considering new laptop
- PIT log issue chooching
- Move script will retain 3(?) zipped logs on machine
- Veeam progress, quote incoming
- PH VPN
- F31 KS ready to go, need to install F31 image on the mountain server
- Office visit
- Drop off Macbook
- Drop off WebIO
- Pick up Mac Mini
- Last week
- OC, Management
- Ao-cravat s/w meeting
- phone screenings
- TMS CoDR GMT
- ESO Esspresso review
- This week
- phone screening
- websockets, AdSec, etc.
- LBC stiched png previewfigure
- max_bytes for ovms
- merge IMS into DMS
- cobbler / kickstart, ims f30
- RPM/CI short tutorial
- TMS CoDR GMT Thu/Fri
- * rm538 broken, monitors rearranged * follow up
- * tradin and LBTC rules * follow up
- Last Week:
- AHS: LBTI alarm patch
- Argos: continued work on RPM build. Attempted to use Heidelberg Ice and TaN RPMs to build, no joy. W. Gaessler reminds me that Argos s/w is particular about lib versions, suggests sticking with source.
- SupportScripts: replace h5csv with dms api call.
- Safety training
- This Week:
- last week
- Second monitor for rm538 broken, monitors rearranged
- Fedora 32 RPM builds
- Fedora install on experimental
- iLocater networking
- This week:
- PIT admin tool backend work. Run into problems with the hot reload (when you make a change to the code you want the browser to reload and show you the changes), not a docker issue but a Django/React issue. There are supposed to be ways to do it but it is not straight forward. One solution involves "ejecting", which means making the webpack code controlling the react project available to the developer to be "hacked". The downside of that is that you can't do upgrades after that, you have to create a new project, "re-eject" again, apply your changes etc or apply the upgrades manually. So for development purposes I'm keeping React and Django apart, each in its own side. Later when moving into operations I'll embed that in Django. Ideally I'll try to have a dev vs build configuration so that doesn't happen manually.
- Safety Training
- Partner with Barry/Dave on testing #IT7952?
- work on configuration of LUCI WebIO, the HTTP GET api seems to be the same but on this device seems to be not consistent. Have to work a bit more on it. David E. provided support for network installation and backup configuration if needed.
- Need to partner with Jenny and revert/test to old GEIRS in LUCI2 osa-luci.sim
- Work on the PIT admin tool backend.
- Partner with Andrew to work on MODS XSLT